You want to have $282,000.00 when you retire 28 years for an annuity. How much money should be deposited at the end of monthly in an investment plan that pays 6.8% compounded monthly, so you will have the $282,000.00 in 28 years?

The monthly payments need to be $. (Round to 2 decimal places.)
Answer #1

Ans $ 281.50

P = Periodic payments
r = rate of interest
n = no of years
Future Value of Annuity = P ( (1 + r)n - 1 ) / r
282000 = P* ((1 + +6.8%/12)^336 - 1) / (6.8%/12)
282000 = 1001.77267779277 * P
P = 282000 / 1001.77267779277
P = 281.50